3a. Consider the line L that passes through the point P0(4, 2, -3) is parallel to the vector


⟶ [ 2

u = -1 <----Matrix

6]


Find a vector equation and the parametric equations of the line L.


b. Find the point of intersection of the line L with the xy-plane (z = 0).



1
Expert's answer
2022-02-23T12:53:44-0500

3.

a.  The vector equation of the line containing "P_0(4, 2, -3)" and parallel to "\vec u=\begin{bmatrix}
 2\\
 -1 \\
6
\end{bmatrix}" is


"\vec r=\begin{bmatrix}
 4+2t\\
2-t \\
-3+6t
\end{bmatrix}"

The corresponding parametric equations are


"x=4+2t, y=2-t, z=-3+6t"

b.

We have the line L


"\dfrac{x-4}{2}=\dfrac{y-2}{-1}=\dfrac{z+3}{6}"

If "z=0"


"\dfrac{x-4}{2}=\dfrac{y-2}{-1}=\dfrac{0+3}{6}"

"x=5, y=3\/2"

Point "(5, 3\/2, 0)" is the point of intersection of the line L with the xy-plane.